### v3

# Value-Free Policy Optimization via Reward Partitioning

Abstract
Single-trajectory reinforcement learning (RL) trains policies from datasets of (prompt, response, reward) triplets, where scalar rewards provide direct supervision. This format aligns with real-world human feedback and avoids the cost of pairwise preference annotations required by methods like Direct Preference Optimization (DPO). Among single-trajectory methods, Direct Reward Optimization (DRO) is effective but depends on value-function approximation, leading to high variance, policy–value coupling, and weak direct supervision.
We propose Reward Partitioning Optimization (RPO), which eliminates value modeling by normalizing rewards via data-driven partitioning. This yields a purely supervised policy objective, free of auxiliary models and joint optimization. RPO thus offers stable, direct supervision and practical simplicity.
Experiments on scalar-feedback language modeling with Flan-T5 show that RPO consistently outperforms single-trajectory baselines including DRO and Kahneman-Tversky Optimization (KTO), establishing RPO as a simple, robust, and theoretically grounded approach to single-trajectory policy optimization.
I Introduction
Aligning large language models (LLMs) typically relies on preference or reward data, via reinforcement learning from human feedback (RLHF) [NEURIPS2023_23e6f78b, dong2024rlhf, ahmadian] or direct preference optimization [rafailov2023direct, tang2024, richemond2024offline, ethayarajh2024]. In RLHF, models are first fine-tuned via supervised learning (SFT) [zhang2023instruction], then a reward model (RM) is trained from human-labeled preference pairs, and finally aligned with reinforcement learning (e.g., PPO [schulman2017proximal], TRPO [schulman2015trust]) to maximize RM scores. This pipeline underlies InstructGPT [ouyang2022training], GPT-4 [achiam2023gpt], and Claude 3 [anthropic2024claude].
RLHF is effective but costly: collecting preference data, training reward models, and RL steps are sample-inefficient and brittle. Reinforcement Learning from AI Feedback (RLAIF) [xu2025, du2023guiding, yuan2023rrhf, wang2025self, yuan2024self] replaces humans with AI evaluators, improving scalability but still depending on large models and introducing inference overhead.
An alternative avoids reward models and RL entirely, directly optimizing policies via supervised learning on preference pairs [rafailov2023direct, tang2024, zhao2023slic]. While simple and stable, high-quality pairwise data is increasingly hard to obtain as LLMs improve, making scaling impractical [grattafiori2024llama, yang2025qwen3, adler2024nemotron].
Single-trajectory learning simplifies supervision: datasets contain (prompt, response, reward) triplets (e.g., thumbs-up/down). Methods like KTO [ethayarajh2024] and DRO [richemond2024offline] train policies directly from these scalar-labeled completions. DRO performs well but requires learning a value function from off-policy data, suffers unstable joint optimization, and supervises the policy only relative to a reference model.
We propose Reward Partitioning Optimization (RPO), which eliminates value estimation by normalizing observed rewards empirically. This yields a direct supervised loss on the policy using only prompt-response-reward triplets and a reference model. RPO is simple, stable, offline-compatible, and scales to large datasets. Our contributions are:
  • Analysis of DRO’s theoretical and practical limitations for single-trajectory learning.
  • Introduction of RPO, replacing value estimation with empirical reward normalization.
  • Validation on instruction-tuning tasks with T5 encoder-decoder models [raffel2020exploring], demonstrating strong performance against DRO and KTO baselines.
II Background
In this section, we review major approaches for aligning large language models (LLMs) with human feedback: reinforcement learning from human feedback (RLHF), direct preference optimization (DPO-style), and direct reward optimization (DRO-style). We highlight their assumptions and limitations, motivating a broader framework that directly leverages scalar reward data.
II-A Reinforcement Learning from Human Feedback (RLHF)
RLHF has become the dominant paradigm for aligning LLMs with human preferences. It proceeds in three main stages:
Supervised Fine-Tuning (SFT): A reference policy $\pi_{\text{ref}}$ is trained using standard supervised learning to mimic high-quality human-written responses.
Reward Model (RM) Training: Given a dataset $\mathcal{D}$ of pairwise preferences $(x,y^{\text{win}},y^{\text{lose}})$—where $x$ is a prompt, and $y^{\text{win}}\succ y^{\text{lose}}$ indicates a preferred response—the goal is to learn a scalar reward function $r_{\phi}(x,y)$ that reflects human preferences. This is typically modeled using a Bradley-Terry [bradley1952rank] formulation:
$\footnotesize\mathbb{P}[y^{\text{win}}\succ y^{\text{lose}}\mid x]=\sigma(r_{\phi}(x,y^{\text{win}})-r_{\phi}(x,y^{\text{lose}})),$ (1)
where $\sigma$ is the logistic sigmoid. The reward model is trained by minimizing the negative log-likelihood:
$\mathcal{L}_{\text{RM}}(r_{\phi})=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}[-\log\sigma(r_{\phi}(x,y^{\text{win}})-r_{\phi}(x,y^{\text{lose}}))]$
Policy Optimization: A policy $\pi_{\theta}$ is then optimized to generate outputs that maximize the learned reward $r_{\phi}(x,y)$ while staying close to the reference policy $\pi_{\text{ref}}$. This is formalized as the following Kullback-Leibler (KL)-regularized objective:
$\footnotesize\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},\,y\sim\pi_{\theta}(\cdot|x)}\left[r_{\phi}(x,y)-\tau\cdot\text{KL}(\pi_{\theta}(y|x)\parallel\pi_{\text{ref}}(y|x))\right],$ (3)
where $\tau>0$ is a regularization coefficient that balances reward maximization and policy deviation. Since this objective is not differentiable with respect to $\pi_{\theta}$, reinforcement learning algorithms such as Proximal Policy Optimization (PPO) [schulman2017proximal] are used to perform the optimization.
Despite its success, RLHF is costly and difficult to implement: reward models require heavy human supervision and risk out-of-distribution errors, while reinforcement learning adds computational overhead and sensitivity to tuning. These challenges motivate alignment methods that avoid explicit RL.
II-B Direct Preference Optimization (DPO-style)
To reduce the cost of full RL, recent methods such as DPO [rafailov2023direct], IPO [azar2024general], and SLiC [zhao2023slic] optimize policies directly from preference pairs without training a reward model or running PPO.
Starting from the KL-regularized objective in Equation 3 and explicitly enforcing the normalization constraint $\sum_{y}\pi_{\theta}(y|x)=1$, the Lagrangian yields a closed-form expression for the optimal policy:
$\pi_{\theta}^{\ast}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\frac{1}{\tau}r_{\phi}(x,y)\right)}{\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\left(\frac{1}{\tau}r_{\phi}(x,y^{\prime})\right)}.$ (4)
Given a dataset of preference pairs
$\mathcal{D}=\{(x_{i},y_{i}^{\text{win}},y_{i}^{\text{lose}})\}_{i=1}^{N},\quad y_{i}^{\text{win}}\succ y_{i}^{\text{lose}},$
we can express the reward difference directly from Equation 4 as
$\footnotesize\frac{r_{\phi}(x_{i},y_{i}^{\text{win}})-r_{\phi}(x_{i},y_{i}^{\text{lose}})}{\tau}=\log\frac{\pi_{\theta}^{\ast}(y_{i}^{\text{win}}|x_{i})/\pi_{\mathrm{ref}}(y_{i}^{\text{win}}|x_{i})}{\pi_{\theta}^{\ast}(y_{i}^{\text{lose}}|x_{i})/\pi_{\mathrm{ref}}(y_{i}^{\text{lose}}|x_{i})}.$ (5)
Minimizing a surrogate loss based on this advantage leads to the general preference-based objective:
$\footnotesize\mathcal{L}_{\text{pref}}(\theta)=\mathbb{E}_{(x,y^{\text{win}},y^{\text{lose}})\sim\mathcal{D}}\Bigg[f\Bigg(\beta\Big(\log\frac{\pi_{\theta}(y^{\text{win}}|x)}{\pi_{\text{ref}}(y^{\text{win}}|x)}-\log\frac{\pi_{\theta}(y^{\text{lose}}|x)}{\pi_{\text{ref}}(y^{\text{lose}}|x)}\Big)\Bigg)\Bigg],$
where $\beta>0$ is a temperature, and $f$ depends on the method:
  • DPO: $f(z)=-\log\sigma(z)$
  • SLiC: $f(z)=\max(0,1-z)$ (hinge loss)
  • IPO: $f(z)=(z-1)^{2}$ (squared loss)
This KL-regularized loss pushes $\pi_{\theta}$ to rank preferred completions above dispreferred ones while staying close to $\pi_{\text{ref}}$, enabling stable training without explicit reward models or RL. However, reliance on curated preference data limits scalability, since real-world feedback often comes as scalar rewards $(x,y,r)$-motivating methods that can exploit such signals directly.
II-C Direct Reward Optimization (DRO-style)
DRO-style methods optimize policies directly from scalar rewards rather than preference pairs. Given a dataset $\mathcal{D}=\{(x_{i},y_{i},r_{i})\}_{i=1}^{N}$, where $r_{i}$ denotes scalar feedback, these methods maximize the utility of desirable responses without explicit reinforcement learning or reward model training.
A notable example is KTO [ethayarajh2024], which draws inspiration from prospect theory [tversky1992advances] to model loss aversion and risk sensitivity. In contrast, DRO [richemond2024offline] starts from the KL-regularized optimal policy in Equation 4:
$\pi_{\theta}^{\ast}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\tfrac{1}{\tau}r(x,y)\right)}{\exp\left(\tfrac{1}{\tau}V^{\ast}(x)\right)},$ (7)
where
$V^{\ast}(x)=\tau\log\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\!\left(\tfrac{1}{\tau}r(x,y^{\prime})\right)$
is the optimal soft value function.
Rearranging Equation 7 yields the consistency condition:
$\footnotesize r(x,y)=V^{\ast}(x)+\tau\log\frac{\pi_{\theta}^{\ast}(y|x)}{\pi_{\mathrm{ref}}(y|x)}.$ (8)
DRO approximates $(\pi^{\ast},V^{\ast})$ by jointly learning a policy $\pi_{\theta}$ and a value function $V_{\phi}$, parameterized by a neural network with parameters $\phi$, through minimization of the Bellman residual:
$\footnotesize\mathcal{L}_{\mathrm{DRO}}(\theta,\phi)=\tfrac{1}{2}\,\mathbb{E}_{(x,y,r)\sim\mathcal{D}}\Big[\big(r-V_{\phi}(x)-\tau\log\tfrac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}\big)^{2}\Big].$ (9)
This objective requires joint optimization of the policy and value network, which can introduce instability due to off-policy value estimation errors. Our proposed RPO eliminates the need for value function learning, simplifying optimization while retaining the advantages of direct reward-based policy learning.
III Method
Direct policy optimization in the RLHF setting with KL regularization commonly relies on the classical exponential form of the optimal policy:
$\footnotesize\pi^{*}_{\theta}(y|x)=\frac{\pi_{\mathrm{ref}}(y|x)\exp\left(\frac{1}{\tau}r(x,y)\right)}{Z^{*}(x)},$ (10)
where the partition function $Z^{*}(x)$ is
$\footnotesize Z^{*}(x)=\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\left(\frac{1}{\tau}r(x,y^{\prime})\right).$ (11)
Current methods such as DRO learn both a policy $\pi_{\theta}$ and a value function $V_{\phi}$ parameterized by $\phi$, where the soft value function relates to the partition function as $V^{*}(x)=\tau\log Z^{*}(x)$. However, this joint learning introduces several difficulties: the value function $V_{\phi}$ can be unstable to approximate, the parameters $(\theta,\phi)$ are tightly coupled, and optimization becomes more complex.
To address these challenges, we propose an alternative method, Reward Partitioning Optimization (RPO), which completely removes the need for an explicit value function model. Instead, RPO performs direct regression on the log-ratio of policies by empirically estimating the partition function, leveraging its analytical form for simpler and more stable training.
Starting from the optimality condition of the policy:
$\footnotesize r(x,y)-V^{*}(x)=\tau\log\frac{\pi^{*}_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)},$ (12)
where the value function $V^{*}(x)$ normalizes the exponentiated rewards, expressible exactly as:
$\footnotesize V^{*}(x)=\tau\log\sum_{y^{\prime}}\pi_{\mathrm{ref}}(y^{\prime}|x)\exp\left(\frac{1}{\tau}r(x,y^{\prime})\right).$ (13)
Instead of learning $V^{*}$ via a separate neural network, RPO estimates it empirically by aggregating samples with the same prompt $x$. Let $\mathcal{I}_{x}=\{i:x_{i}=x\}$ index all samples for prompt $x$. Then, the empirical partition function and value are:
$\footnotesize\widehat{Z}(x)=\sum_{j\in\mathcal{I}_{x}}\pi_{\mathrm{ref}}(y_{j}|x)\exp\left(\frac{1}{\tau}r(x,y_{j})\right),\quad\widehat{V}(x)=\tau\log\widehat{Z}(x).$ (14)
RPO minimizes the squared error between the log-policy ratio and the scaled centered reward, optimizing only over the policy parameters $\theta$. Formally, the loss can be expressed as the expectation:
$\footnotesize\mathcal{L}_{\mathrm{RPO}}(\theta)=\frac{1}{2}\mathbb{E}_{(x,y,r)\sim\mathcal{D}}\left[\left(\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}-\frac{1}{\tau}(r(x,y)-V^{*}(x))\right)^{2}\right].$ (15)
In practice, we approximate this with the empirical sum over the training set $\{(x_{i},y_{i},r_{i})\}_{i=1}^{N}$ and use $\widehat{V}(x)$ as an estimate for $V^{*}(x)$:
$\footnotesize\hat{\mathcal{L}}_{\mathrm{RPO}}(\theta)=\frac{1}{2N}\sum_{i=1}^{N}\left(\log\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\mathrm{ref}}(y_{i}|x_{i})}-\frac{1}{\tau}\left(r_{i}-\widehat{V}(x_{i})\right)\right)^{2}.$ (16)
Defining the residual for sample $i$ as:
$\footnotesize\delta_{i}(\theta):=\log\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\mathrm{ref}}(y_{i}|x_{i})}-\frac{1}{\tau}\left(r_{i}-\widehat{V}(x_{i})\right),$ (17)
we compute the gradient of the loss with respect to the policy logits $z_{\theta}(y|x)$, where the policy is parameterized via a softmax:
$\footnotesize\pi_{\theta}(y|x)=\frac{\exp(z_{\theta}(y|x))}{\sum_{y^{\prime}}\exp(z_{\theta}(y^{\prime}|x))}.$ (18)
The gradient of the loss is then:
$\footnotesize\nabla_{z_{\theta}(y|x)}\hat{\mathcal{L}}_{\mathrm{RPO}}(\theta)=\frac{1}{N}\sum_{i:x_{i}=x}\delta_{i}(\theta)\cdot\nabla_{z_{\theta}(y|x)}\log\pi_{\theta}(y_{i}|x).$ (19)
Using the standard softmax gradient, we have:
$\footnotesize\nabla_{z_{\theta}(y|x)}\log\pi_{\theta}(y_{i}|x)=\mathbf{1}_{y=y_{i}}-\pi_{\theta}(y|x),$ (20)
where $\mathbf{1}_{y=y_{i}}$ is the indicator function.
Substituting into Equation (19), the final gradient expression becomes:
$\footnotesize\nabla_{z_{\theta}(y|x)}\hat{\mathcal{L}}_{\mathrm{RPO}}(\theta)=\frac{1}{N}\sum_{i:x_{i}=x}\delta_{i}(\theta)\left(\mathbf{1}_{y=y_{i}}-\pi_{\theta}(y|x)\right).$ (21)
This gradient has an intuitive interpretation as the weighted discrepancy between the one-hot target and the policy prediction, weighted by the residual error $\delta_{i}(\theta)$. Crucially, RPO avoids backpropagation through the value estimate $\widehat{V}(x)$, resulting in a simpler, more stable training procedure that retains the theoretical benefits of DRO without learning an explicit value function.
1
Input: Dataset $\mathcal{D}=\{(x_{i},y_{i},r_{i})\}_{i=1}^{N}$, initial policy $\pi_{\theta}$, reference policy $\pi_{\mathrm{ref}}$, temperature $\tau>0$, number of epochs $T$, learning rate $\eta$
Output: Trained policy $\pi_{\theta}$
2
3Group indices by prompt: $\forall x,\mathcal{I}_{x}=\{i:x_{i}=x\}$;
4
5for $t\leftarrow 1$ to $T$ do
6 foreach batch $B$ from $\mathcal{D}$ do
7    foreach $(x_{i},y_{i},r_{i})\in B$ do
8         Compute $\ell_{i}\leftarrow\log\pi_{\theta}(y_{i}|x_{i})$;
9         Compute $\ell_{i}^{\mathrm{ref}}\leftarrow\log\pi_{\mathrm{ref}}(y_{i}|x_{i})$;
10
11        Estimate empirical partition: $\widehat{Z}(x_{i})=\sum_{j\in\mathcal{I}_{x_{i}}}\pi_{\mathrm{ref}}(y_{j}|x_{i})\exp\left(\frac{1}{\tau}r(x_{i},y_{j})\right)$;
12
13        Compute $\widehat{V}(x_{i})\leftarrow\tau\log\widehat{Z}(x_{i})$;
14         Compute $t_{i}\leftarrow\frac{r_{i}-\widehat{V}(x_{i})}{\tau}$;
15         Compute $\ell_{\mathrm{RPO},i}\leftarrow\frac{1}{2}\left(\ell_{i}-\ell_{i}^{\mathrm{ref}}-t_{i}\right)^{2}$;
16
17     Update $\theta$ by gradient descent on $\frac{1}{|B|}\sum_{i\in B}\ell_{\mathrm{RPO},i}$;
18
19
20return $\pi_{\theta}$;
Algorithm 1 RPO Optimization
To operationalize the RPO training procedure described above, we present a concrete algorithmic implementation (Algorithm 1). The algorithm iteratively estimates the target log-ratio using a partition function approximation based on the reference policy, computes the residual loss for each sample, and updates the policy parameters via gradient descent.
Property DRO RPO
Explicit $V$ estimation Yes (model $V_{\phi}$) No (empirical estimation)
[…]
Optimization complexity Medium to high Simpler, more stable
Robustness to estimation errors Sensitive Less sensitive
Practical deployment Requires tuning More direct, no $V$ tuning
TABLE I: Comparison between DRO and RPO methods.
To further highlight the distinctions between the proposed RPO method and standard distributionally robust optimization (DRO) approaches, Table I summarizes key differences in terms of implementation, coupling, and practical considerations.
The RPO method formulates policy learning as a direct regression on the log-ratio between target and reference policies, using rewards and an empirical partition estimation. This removes the need to train a value function, simplifying the learning procedure and enhancing numerical stability.
IV Experiments
IV-A Dataset and Base Models
We follow the experimental setup of Richemond et al. [richemond2024offline] and use the UltraFeedback dataset [cui2023ultrafeedback], which consists of offline triplets $(\text{prompt},\text{completion},\text{reward})$ with multiple completions per prompt. Rewards are standardized to zero mean and unit variance for training stability.
We use instruction-tuned Flan-T5 models [roberts2023scaling] at three scales: Small (77M), Large (783M), and XL (2.85B). All models are initialized from FLAN instruction-tuned checkpoints [chung2024scaling], which serve as the reference policy $\pi_{\text{ref}}$, enabling fair comparison across model sizes.
IV-B Evaluation
We evaluate models on a validation set of 1,000 high-quality examples (reward =10.0=10.0) from the UltraFeedback dataset [cui2023ultrafeedback]. This subset is used for two complementary evaluation protocols.
Automatic Metrics. During training, we monitor performance using standard generation metrics: BLEU, ROUGE-L, METEOR, BERTScore, SacreBLEU, and Toxicity. These metrics capture surface-level similarity and general quality trends but are limited in assessing alignment and preference satisfaction.
LLM-Based Preference Evaluation. To obtain more human-aligned judgments, we perform post-training pairwise evaluations using the LLM-as-a-judge paradigm [gu2024survey]. Model outputs are compared using Mistral v3 [jiang2023mistral7b] and LLaMA v3 [touvron2023llamaopenefficientfoundation] as impartial judges. Given an instruction and two responses, the judge selects the better one based on helpfulness and relevance, or Neutral if neither is clearly preferred.
Allowing neutral judgments mitigates bias from forced binary decisions and enables more reliable preference-sensitive evaluation in RLHF settings.
IV-C Training Setup and Baseline Comparison
All models are trained with the AdamW optimizer [loshchilov2019decoupled] using a learning rate of $1\times 10^{-4}$. We apply a linear warmup and decay schedule with 150 warmup steps, where the total number of training steps is given by
$\texttt{num\_training\_steps}=\texttt{steps\_per\_epoch}\times\texttt{epochs}.$
We compare RPO against KTO and DRO baselines. To isolate methodological differences, the temperature parameter is fixed to $\tau=1.0$ across all experiments.
Figure 1: Average training time in hours for KTO, DRO, and RPO across FLAN-T5 model sizes: Small (77M parameters), Large (783M parameters), and XL (2.85B parameters). RPO consistently converges faster, demonstrating superior training efficiency.
Figure 1 reports average training time across FLAN-T5 Small, Large, and XL models. RPO consistently converges faster than KTO and DRO at all scales, demonstrating superior training efficiency and scalability.
IV-D Automatic Metric-Based Validation
We evaluate model quality during training using standard automatic generation metrics across three methods (KTO, DRO, and RPO) and three model sizes (Small, Large, XL). The metrics include BERTScoreF1, ROUGE-L, BLEU, SacreBLEU, METEOR, and average toxicity, where $\uparrow$ indicates higher-is-better and $\downarrow$ lower-is-better.
Metric KTO DRO RPO
Small Large XL Small Large XL Small Large XL
BERTScoreF1 0.8499 0.8401 0.7883 0.8562 0.8591 0.8537 0.8661 0.8706 0.8702
ROUGE-L 0.2369 0.2491 0.2480 0.2522 0.3034 0.2502 0.2885 0.3162 0.2969
BLEU 0.0293 0.0438 0.0108 0.0173 0.0274 0.0419 0.0360 0.0464 0.0511
SacreBLEU 2.8740 4.3785 1.3653 1.7218 2.8589 4.2672 3.6013 4.7903 5.1063
METEOR 0.2083 0.2152 0.2030 0.2029 0.2337 0.2153 0.2418 0.2632 0.2567
Toxicityavg 0.0133 0.0241 0.0242 0.0146 0.0130 0.0122 0.0125 0.0093 0.0078
TABLE II: Automatic metric-based validation results for KTO, DRO, and RPO across different model sizes.
Table II shows that RPO consistently outperforms KTO and DRO across semantic similarity and fluency metrics while achieving lower toxicity at all model scales. These results indicate stronger preference alignment, though automatic metrics alone are insufficient to fully capture human judgments. We therefore complement this analysis with LLM-based preference evaluation in the following subsection.
IV-E LLM-Based Preference Evaluation
We evaluate preference alignment using two LLM judges, Mistral and LLaMA, comparing RPO against SFT, KTO, and DRO across Small, Large, and XL models. We report the win rate (%) of RPO over each baseline, with standard deviation.
Model Size Mistral Judge LLaMA Judge
SFT KTO DRO SFT KTO DRO
Small 82.78 $\pm$ 1.93 82.60 $\pm$ 2.39 83.97 $\pm$ 2.34 63.66 $\pm$ 1.62 65.76 $\pm$ 0.59 65.26 $\pm$ 0.41
Large 84.20 $\pm$ 1.64 86.53 $\pm$ 1.77 84.86 $\pm$ 2.05 73.02 $\pm$ 0.52 72.30 $\pm$ 1.72 74.81 $\pm$ 0.63
XL 83.01 $\pm$ 2.39 93.62 $\pm$ 5.88 83.19 $\pm$ 0.88 68.43 $\pm$ 1.17 61.18 $\pm$ 9.33 67.15 $\pm$ 3.60
TABLE III: Win rate (%) of RPO over SFT, KTO, and DRO as judged by Mistral and LLaMA. Higher is better.
Table III shows that RPO consistently outperforms all baselines across model sizes and judges, with particularly strong margins under the more discriminative Mistral judge. These results indicate that RPO produces responses better aligned with human preferences than competing methods.
RPO’s advantage stems from directly optimizing the policy via preference ranking, avoiding explicit value modeling required by KTO and DRO. This simpler optimization leads to more stable training and improved preference alignment across scales.
IV-F Impact of Regularisation Parameter $\tau$
We study the effect of the temperature parameter $\tau$ in RPO by comparing its win rate over SFT for $\tau\in\{0.1,1.0,5.0\}$ using Mistral and LLaMA judges. Results for $\tau=1.0$ match those reported in Section IV-E.
Model Size Mistral Judge LLaMA Judge
$\tau=0.1$ $\tau=1.0$ $\tau=5.0$ $\tau=0.1$ $\tau=1.0$ $\tau=5.0$
Small 84.45 $\pm$ 1.14 82.78 $\pm$ 1.93 83.03 $\pm$ 1.55 78.73 $\pm$ 1.07 63.66 $\pm$ 1.62 66.52 $\pm$ 1.55
Large 59.50 $\pm$ 19.62 84.20 $\pm$ 1.64 83.72 $\pm$ 1.05 56.40 $\pm$ 12.21 73.02 $\pm$ 0.52 74.86 $\pm$ 1.63
XL 90.24 $\pm$ 6.66 83.01 $\pm$ 2.39 82.81 $\pm$ 3.76 82.07 $\pm$ 5.30 68.43 $\pm$ 1.17 66.28 $\pm$ 2.48
TABLE IV: Win rate (%) of RPO over SFT for different values of $\tau$, evaluated by Mistral and LLaMA.
Table IV shows that RPO is robust to the choice of $\tau$, with strong and stable performance at $\tau=1.0$ and $\tau=5.0$ across model sizes and judges. Very low temperature ($\tau=0.1$) can introduce instability at larger scales, particularly under Mistral.
Overall, these results indicate that RPO does not require precise temperature tuning, supporting its practicality and stability compared to methods that rely on explicit value modeling or tightly coupled reward calibration.
IV-G Ablation Study: Importance of Empirical Normalization
We study the role of empirical normalization in RPO by comparing the full method against an ablated variant, RPO-noNorm, which removes the baseline $\widehat{V}(x)$ and directly optimizes raw rewards. We evaluate both methods across $\tau\in\{0.1,1.0,5.0\}$ and model sizes (Small, Large, XL).
Automatic Metrics. Table V shows that RPO consistently outperforms RPO-noNorm across all metrics, temperatures, and model scales. RPO remains stable under temperature variation and benefits from scaling, whereas RPO-noNorm exhibits degraded fluency, weaker semantic alignment, higher toxicity, and minimal gains from larger models.
RPO RPO-noNorm
$\tau=0.1$ $\tau=1.0$ $\tau=5.0$ $\tau=0.1$ $\tau=1.0$ $\tau=5.0$
Metric S L XL S L XL S L XL S L XL S L XL S L XL
BERTScoreF1 0.8648 0.8702 0.8691 0.8661 0.8706 0.8702 0.8653 0.8697 0.8684 0.7734 0.7681 0.7702 0.7741 0.7689 0.7712 0.7726 0.7705 0.7723
ROUGE-L 0.2861 0.3143 0.2951 0.2885 0.3162 0.2969 0.2872 0.3150 0.2958 0.2063 0.2028 0.2051 0.2089 0.2043 0.2067 0.2115 0.2078 0.2101
BLEU 0.0354 0.0460 0.0503 0.0360 0.0464 0.0511 0.0348 0.0456 0.0498 0.0130 0.0127 0.0139 0.0135 0.0129 0.0143 0.0131 0.0132 0.0137
SacreBLEU 3.5421 4.7650 5.0228 3.6013 4.7903 5.1063 3.5470 4.7582 5.0157 1.2360 1.2203 1.2632 1.2743 1.2312 1.2914 1.2634 1.2398 1.2665
METEOR 0.2401 0.2619 0.2551 0.2418 0.2632 0.2567 0.2403 0.2626 0.2559 0.1749 0.1694 0.1726 0.1762 0.1708 0.1736 0.1780 0.1719 0.1741
Toxicityavg 0.0129 0.0095 0.0079 0.0125 0.0093 0.0078 0.0130 0.0097 0.0081 0.0243 0.0250 0.0247 0.0245 0.0251 0.0249 0.0239 0.0244 0.0242
TABLE V: Automatic metric comparison of RPO and RPO-noNorm across temperatures and model sizes.
LLM-Based Preference Evaluation. As shown in Table VI, RPO is preferred over RPO-noNorm in over 90% of comparisons across all temperatures and model sizes, under both Mistral and LLaMA judges.
Model Size Mistral LLaMA
$\tau=0.1$ $\tau=1.0$ $\tau=5.0$ $\tau=0.1$ $\tau=1.0$ $\tau=5.0$
Small 96.22 $\pm$ 0.91 95.04 $\pm$ 0.84 95.77 $\pm$ 1.02 92.75 $\pm$ 1.13 91.34 $\pm$ 1.25 91.80 $\pm$ 0.97
Large 94.85 $\pm$ 0.77 96.13 $\pm$ 0.66 95.90 $\pm$ 0.58 90.64 $\pm$ 1.31 91.76 $\pm$ 0.89 92.08 $\pm$ 0.74
XL 97.34 $\pm$ 0.52 95.81 $\pm$ 0.63 96.12 $\pm$ 0.49 94.21 $\pm$ 1.03 92.99 $\pm$ 0.85 93.17 $\pm$ 0.77
TABLE VI: LLM preference win rate (%) of RPO over RPO-noNorm.
These results demonstrate that empirical normalization via $\widehat{V}(x)$ is critical for stable and effective optimization. Temperature tuning alone cannot compensate for its removal, confirming normalization as a key component of RPO.
V Qualitative Study of Prompt Completions
We present qualitative examples to illustrate how different optimization strategies affect response quality on real-world prompts. We focus on cases where both LLaMA and Mistral judges prefer completions generated by RPO over DRO and KTO. Across these examples, RPO produces responses that are more complete, informative, and well-structured, reflecting stronger alignment with human preferences and reward signals. In contrast, DRO and KTO generate reasonable but comparatively shallow answers, often lacking nuance or contextual depth. All examples are generated using the XL model variant, where differences between methods are most pronounced.
Prompt 1: “How can managers build trust with remote teams while maintaining accountability?”
RPO: Managers should build trust through consistent communication, transparency about goals, and empathy toward individual circumstances. Regular check-ins that focus on support rather than surveillance help establish accountability without eroding trust.
DRO: Build trust with open communication. Hold people accountable by setting deadlines.
KTO: Use video calls and check-ins to build trust and keep track of work.
Prompt 2: “Explain why excessive screen time may negatively affect children’s cognitive development.”
RPO: Excessive screen time in children is associated with reduced attention span, impaired sleep, and lower engagement in cognitively stimulating activities such as reading or creative play. These factors may contribute to slower cognitive and emotional development.
DRO: Too much screen time can lead to worse focus and bad sleep habits in children.
KTO: It may reduce focus and sleep.
Prompt 3: “Is nuclear energy a viable solution for climate change mitigation?”
RPO: Nuclear energy offers a low-carbon, high-output alternative to fossil fuels and can significantly reduce greenhouse gas emissions. While it presents challenges such as waste disposal and high initial costs, it remains a viable component of a broader climate strategy.
DRO: Nuclear energy is low-carbon but has risks like waste and cost.
KTO: It helps reduce emissions but is expensive.
TABLE VII: Qualitative comparison of prompt completions generated by the XL model. RPO outputs are preferred by both LLaMA and Mistral judges and are consistently more informative and coherent than those produced by DRO and KTO.
VI Conclusion
We introduced RPO, a simple and scalable fine-tuning method that normalizes observed rewards to eliminate reward models and RL loops. RPO directly supervises policies against reward-labeled completions while staying close to a reference, achieving strong performance over SFT, DRO, and KTO across model sizes. It is easy to implement, stable, and robust to hyperparameter variation, as confirmed by both Mistral and LLaMA evaluations.
While effective, RPO relies on the quality and coverage of observed data and the alignment of the reference policy, which may limit generalization beyond the dataset.
Future Work. We plan to scale RPO to larger GPT-style models, evaluate on broader datasets including human and multi-turn feedback, and explore dynamic references or lightweight value learning to improve coverage without added complexity.